OpenStudy (anonymous):

Factor. Note that the middle term has already been split. a^2+5a+6a+30 Factor. Note that the middle term has already been split. a^2+5a+6a+30 @Mathematics

OpenStudy (anonymous):

makes it very easy (a + 5)(a + 6)

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

joyce153 - factoring requires understanding the expanded for of (x+a)(x+b) = x^2+ax+bx+ab = x^2 + (a+b)x + ab. So, anytime you are asked to factor, you need to find two numbers that add up to the coefficient of x such that the product of those two numbers is equal to the constant term in that expression.